Home - Search - Browse - Alphabetic Index: 0- 1- 2- 3- 4- 5- 6- 7- 8- 9
A- B- C- D- E- F- G- H- I- J- K- L- M- N- O- P- Q- R- S- T- U- V- W- X- Y- Z
More Details for 2008-02-29
ISS On-Orbit Status 02/29/08

Today is February's fifth Friday, because of its 29 Leap Year days; the last time February had 5 Fridays was in 1980 and next time will be in 2036.

Before breakfast and exercise, FE-2 Eyharts performed his first PHS (Periodic Health Status) w/Blood Labs examination. CDR Whitson assisted in drawing blood and using the U.S. PCBA(Portable Clinical Blood Analyzer). The second part of PHS, Subjective Clinical Evaluation, was performed later in the day. (The PHS exam, with PCBA analysis and clinical evaluation, is guided by special software (IFEP, In-Flight Examination Program) on the MEC (Medical Equipment Computer). While PCBA analyzes total blood composition, the blood's hematocrit is particularly measured by the Russian MO-10 protocol.)

In preparation for 1J/A arrival and transfer of the JAXA ELM-PS (Experiment Logistics Module - Pressurized Section), called JLP for short, the CDR installed the CBCS (Centerline Berthing Camera System) at the Node-2 zenith (upper) hatch, powered it up and checked it out. Afterwards, she disconnected the CBCS electronics cables to avoid hatch 'dragthroughs'. (JLP will be transferred from the Shuttle cargo bay with the SSRMS on 3/14 and berthed at the Node-2 zenith CBM (Common Berthing Mechanism) during EVA-1 (of five EVAs total) by Rick Linnehan (EV1) & Garrett Reisman (EV2).)

Afterwards, Whitson moved to the Airlock (A/L), connected EMU (Extravehicular Mobility Unit) #3008 to the SCU (Service & Cooling Umbilical) and initiated the standard 1-hr scrubbing process on the spacesuit's cooling water loops, filtering ionic and particulate matter (via a 3-micron filter). No iodination for biocide maintenance was required this time. The cooling loops were afterwards reconfigured and the EMU water processing kit disassembled and stowed. (This prepared EMU #3008 for use on STS-123/docked.)

FE-1 Malenchenko had 4 hrs set aside for a major IFM (Inflight Maintenance) on the Russian Thermal Control System (STR), to restore its Loop 1 (KOB1) by removing two units (BS) containing electric recirculation pumps (ENA) on the 3SPN2 pump panel and replacing them with spares from FGB stowage. (Such R&Rs have been done before by previous crews.)

Later, Malenchenko transferred the KAR01 SIS-1 kit from the DC1 and set it up in the Service Module (SM) for exposure and photography in preparation for the next SFP (Spaceflight Participant), arriving on Soyuz TMA-12/16S in April. (KAR01 SIS-1, part of the KAR experiment suite of SFP/Visiting Cosmonaut Ko San from South Korea, a researcher at Samsung Advanced Institute of Technology, studies the growth pattern of vegetable cultures like soy, rice, radish, orchid seeds and others in the zero-G environment.)

In the COL (Columbus Orbital Laboratory), after yesterday's preparatory work on the FSL (Fluid Science Laboratory) facility, Leo Eyharts today set up and configured the CEM-L (Central Experiment Module-Lower) for its activation and initiation of the first experiment run on Monday (3/3). (ESA will attempt to complete five experiment runs within FSL prior to Flight 1J/A.).

Peggy Whitson powered up the MSG (Microgravity Science Glovebox) and performed another session with the InSPACE-2 (Investigating the Structure of Paramagnetic Aggregates from Colloidal Emulsions) experiment, today conducting runs #36 & #37, with a vial change from #002 to #007 between the runs and an increase in frequency from 0.66 Hz to 2Hz. Afterwards, the MSG was deactivated again. (InSPACE obtains basic data on MR (magnetorheological) fluids, i.e., a new class of "smart materials" that can be used to improve or develop new brake systems, seat suspensions robotics, clutches, airplane landing gear, and vibration damper systems. The colloidal (dispersed) particles are contained in CAs (Coil Assemblies) in the MSG that subject them to electric fields at certain strength and frequencies. The desired strong dipolar interaction between the small colloidal particles can be achieved in micro-G simply with an external magnetic field being turned on and off. On the ground, the flow properties (rheology) of many materials, especially those making up consumer products like detergents, fabric softeners, toothpaste and paints, are similarly controlled, though not by magnetic fields but by adding a polymer. It now appears, for example, that new formulations of fabric softeners may perform better in space than on earth.)

Yuri Malenchenko worked through a long list of standard maintenance/service tasks on Russian Segment (RS) systems. Specifically, the FE-1 -

Completed the periodic (generally monthly) service of the ESA/RSC-Energia experiment ALTCRISS (Alteino Long Term monitoring of Cosmic Rays on the ISS), removing its PCMCIA memory card (#943) from the AST spectrometer (inserted on 2/15) and replacing it with a new card (#944), then, as usual, checking the used card on the RSK1 laptop for quantity and size of files;
Employed the Russian IPD Draeger tube and CMS (Countermeasure System) samplers to take air samples for Carbon Monoxide (CO), to verify data obtained with the IPD-CO sampler on 2/25;
Took readings of potentially harmful contaminants in the SM with the CMS (part of the GANK-4M Real-Time Harmful Contaminant Gas Analyzer system), which uses preprogrammed microchips to measure Formaldehyde (H2CO, methanal), CO and Ammonia (NH3), taking one measurement per microchip;
Performed monthly maintenance on the deactivated Russian IK0501 GA (Gas Analyzer) of the SOGS Pressure Control & Atmospheric Monitoring System by replacing its CO2 filter assembly (BF) with a new unit from FGB stowage (done last: 1./21), then reactivating the unit;
Completed the routine servicing of the SOZh system (Environment Control & Life Support System, ECLSS) in the SM (Regular daily SOZh maintenance consists, among else, of checking the ASU toilet facilities, replacement of the KTO & KBO solid waste containers and replacement of EDV-SV waste water and EDV-U urine containers);
Linked up with TsUP-Moscow stowage specialists at ~4:25am EST via S-band to conduct the weekly IMS (Inventory Management System) tagup, discussing stowage issues and equipment locations;
Conducted the daily IMS maintenance, updating/editing its standard 'delta file' including stowage locations, for the regular weekly automated export/import to its three databases on the ground (Houston, Moscow, Baikonur); and - working from the Russian voluntary 'time permitting' task list -
Completed the regular daily checkup on the Japanese experiment GCF-JAXA (Granada Crystallization Facility) in the Russian TBU incubator, maintained at +20 degC, including a temperature check on its ART (automatic temperature recorder), and
Performed the daily monitoring, picture-taking and downloading on the BIO-5 Rasteniya-2 ("Plants-2") experiment which researches growth and development of plants (peas) under spaceflight conditions in the Lada-12 greenhouse from IBMP (Institute of Bio-Medical Problems, Russian: IMBP).
Peggy gathered measurements for the weekly atmospheric status check for ppO2 (Partial Pressure Oxygen) and ppCO2 (pp Carbon Dioxide), using the hand-held CSA-CP (Compound Specific Analyzer-Combustion Products), CSA-O2 (CSA -Oxygen sensor) and CDMK (CO2 Monitoring Kit). Batteries were to be replaced if necessary. (Purpose of the 15-min activity is to trend with MCA (Major Constituents Analyzer), i.e., to correlate the hand-held readings with MCA measurements. CSA-CP sensors (and readings) employed in the SM were #1051 (20.1%) & #1044/prime (20.3%); in Node-1 #1058 (19.0%); and in Node-2 #1045 (20.0%). O2 sensor checks used #1052 (20.8%), #1041 (20.9%). CDMK CO2 level in Lab was 0.5% and in SM 0.5%.)

As a fix for the erratic temperature sensor A of the CDRA (Carbon Dioxide Removal Assembly)'s absorbent bed #2, Whitson today installed a jumper cable to bypass the sensor, then supported the ground-commanded activation of CDRA (9:50am-11:50am) by connecting the regular ITCS LTL (Internal Thermal Control System/Low Temperature Loop) coolant jumper at the LAB1D6 rack for cooling. (The anomalous dP signature has not been explained yet, despite several days of troubleshooting.)

Whitson and Eyharts conducted a 30-min review of new POC DOUG (Portable Onboard Computer/Dynamic Onboard Ubiquitous Graphics) material, now using the new R5 software upgrade which mainly incorporates SPDM functionality into the MSS (Mobile Service System) and PCS (Portable Computer System) displays. The review was for the SSRMS (Space Station Remote Manipulator System) activities scheduled on 3/4 (Tuesday). (Next week's Robotics ops will essentially be the regular prelaunch checkout of the MSS/SSRMS complex for STS-139/1J/A, involving maneuvers to grapple the MBS PDGF-3 (Mobile Base System Power & Data Grapple Fixture 3), grapple & ungrapple the Node-2 PDGF, perform a checkout of the Cupola & Lab RWS (Robotics Workstation), and finish by moving the robotarm to the 1J/A start position.)

Later today, before sleep time, Peggy will stand by to reconnect PCS laptops after ground specialists have finished uploading patches to the three C&C MDMs (Command & Control Multiplexer/Demultiplexer) computers. (For loading the software patches (which deal with ATV Fire Warning and the Time Tag Queue), engineers have performed two Backup vs.Standby C&C swaps, followed by a Primary-to-Standby C&C transition about 12 hrs later. The CDR will reconnect PCSs after the C&C1 transition to Primary (C&C3 then being Backup & C&C2 Standby).)

Leo Eyharts had another 60 minutes for himself for general orientation (station familiarization & acclimatization), as is standard daily rule for the first two weeks after starting station residency.

The crewmembers completed their regular 2.5-hr. physical workout program (about half of which is used for setup & post-exercise personal hygiene) on the CEVIS cycle ergometer (CDR, FE-2), TVIS treadmill (FE-1), RED resistive exercise device (CDR, FE-2) and VELO bike with bungee cord load trainer (FE-1, FE-2).

Peggy then transferred the crew's exercise data file to the MEC laptop for downlinking, as well as the daily wristband HRM (Heart Rate Monitor) data of the workouts on RED, followed by their erasure on the HRM storage medium (done six times a week).

At ~3:45am EST, the crew held the regular (nominally weekly) tagup with the Russian Flight Control Team (GOGU), including Shift Flight Director (SRP), at TsUP via S-band/audio, phone-patched from Houston and Moscow.

At ~4:45am, Leo conducted the regular weekly 15-min. audio tagup with ESA's Col-CC (Columbus Control Center) in Oberpfaffenhofen near Munich/Germany. (S/G-2 (Space-to-Ground 2) phone patch via SSC-10 (Station Support Computer 10)).

At ~2:10pm, the crew had their twelfth weekly tagup with the Lead Flight Director at JSC/MCC-H via S-band/audio. (S/G-2 (Space-to-Ground 2) phone patch via SSC-10 (Station Support Computer 10)).

OGS Activation Update: By adding some water into the US OGS (Oxygen Generation System) WDS (Water Delivery System), the ground succeeded last night in activating the OGS. The system is currently generating O2 at a nominal 3-person equivalent rate.

COL Update: Columbus module and systems are performing nominally.

CWC Update: An updated CWC (Contingency Water Container) 'cue card' was uplinked for the crew's reference. (The new card (16-0018V) lists 35 CWCs (~1427.9 L total) for the four types of water identified on board: technical water (812.9 L, for Elektron, flushing, hygiene), potable water (559.6 L), condensate water (54.7 L), waste/EMU dump and other (07 L). Of the 35 containers, nine CWCs with technical water (388.9 L) cannot be used until cleared for Wautersia bacteria, and 10 CWCs with potable water (427 L), transferred from Atlantis, are not cleared for use pending analysis of samples returning on 1E.)

CEO photo targets uplinked for today were Hubbard Brook, New Hampshire (there was snow on the ground and ice on the rivers this time of year for the first of two Long Term Ecological Research (LTER) sites. The main objective from the standpoint of CEO is a record of land cover/land use change on a seasonal basis. The Hubbard Brook Experimental Forest (HBEF) is a 3,160 hectare reserve located in the White Mountain National Forest, near Woodstock, New Hampshire. The on-site research program is dedicated to the long-term study of forest and associated aquatic ecosystems. Overlapping imagery of this area was requested), and Plum Island Ecosystem, Maine (the second of two LTER sites for this day, the principal study site was the Plum Island Sound estuary, its coupled Parker, Rowley and Ipswich River watersheds and the adjacent coastal ocean, the Gulf of Maine. This LTER site focuses on how several aspects of global change influence organic matter and nutrients from land, ocean and marshes and how they interact with the external drivers (climate, land use, river discharge, sea level). Overlapping imagery of this area was requested).

Home - Search - Browse - Alphabetic Index: 0- 1- 2- 3- 4- 5- 6- 7- 8- 9
A- B- C- D- E- F- G- H- I- J- K- L- M- N- O- P- Q- R- S- T- U- V- W- X- Y- Z
© 1997-2017 Mark Wade - Contact
© / Conditions for Use